OpenStudy (anonymous):

The equation, y = 10(1.0605)^x, represents the winning prize money, in thousands, for first place in the Indianapolis 500 auto race. x represents the number of years after 1911, the first year. What is the value for 1911 (x = 0)? This is the prize money, in thousands. For instance, if you got 23, this represents $23,000. A. $10,605 B. $10,000 C. $0

OpenStudy (anonymous):

10,000 y = 10 (1.0605)^0 = 10*1 = 10,000

OpenStudy (anonymous):

y = 10(1.0605)^x, where x=0 y = 10(1.0605)^0 = 10(1) = 10 B. $10,000